An employer interviews 8 people for 4 technical support positions in a company, and 3 of the 8 people are men. If all 8 are qualified, in how many ways can the employer fill the technical support positions if exactly 2 are men
Found 2 solutions by ewatrrr, stanbon:
Answer by ewatrrr(24785)

 
Hi,
An employer interviews 8 people for 4 technical support positions in a company, and 3 of the 8 people are men.
If all 8 are qualified, in how many ways can the employer fill the technical support positions if exactly 2 are men
3C2*5C2= 3*10 = 30 ways

Answer by stanbon(75887)

An employer interviews 8 people for 4 technical support positions in a company, and 3 of the 8 people are men. If all 8 are qualified, in how many ways can the employer fill the technical support positions if exactly 2 are men
-----
# of ways to pick 2 men: 3C2 = 3
# of ways to pick 2 women: 5C2 = 10
----
# of ways to satisfy the problem condition: 3*10 = 30
=========================================================
Cheers,
Stan H.
==============